OpenStudy (anonymous):

One day in march in Jilton Head ,South Caolina, the first high tide occurred at 6:18am.THe hight tide was 7.05 feet ,and the low tide was -0.30 feet .The period for the oscillation of the tides is 12 hours and 24 minutes .

OpenStudy (anonymous):

a) Determine what time the next hight will occur b) Write the period of the oscillation as a decimal c) What is the amplitude of the sinusoidal function that models the tide? d) If t=0 represents midnight, write a sinusoidal function that models the tide. e) At what time will the tides be at 6 feet for the first time that day?

OpenStudy (anonymous):

my answer is a) amplitude = (7.05-0.3)/2=3.675 period = 2pi/12.24=6.24 6.18+12.24=6 hours 42 minutes b) = 12.4 hours \[c) \pm 3.675 feet\] i can get answer a,b,c, I need help d and e only

OpenStudy (anonymous):

the key for d) is \[y= 3.375+3.675 \sin( \frac{\pi }{6.2}t-1.62)\] d) is 4:46 AM